OpenStudy (anonymous):

how do i find 125^2/3

OpenStudy (raden):

first, 125 = 5^3 so, we have the equation (5^3)^(2/3) then use the property of exponent : (a^m)^n = a^(m * n) so, (5^3)^(2/3) = 5^(3 * 2/3) = 5^2 = 25

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."